An electronics store received a shipment of 20 calculators, including 7 defective. 4 of the calculators are selected to be sent to a high school.
A)How many selections can be made using the original shipment?
B) How many selections will contain defective calculators?
C) How many will contain 1 defective calculator?
D) How may of these will contain 3 defective calculators?

Answer by stanbon(75887)

An electronics store received a shipment of 20 calculators, including 7 defective. 4 of the calculators are selected to be sent to a high school.
A)How many selections can be made using the original shipment?::
20C4 = (20*19*18*17)/(1*2*3*4) = (5*19*3*17) = 4845
--------------------------------------------------
B) How many selections will contain defective calculators?
# with NO defective = 13C4 = (13*12*11*10)/(1*2*3*4) = (13*11*5) = 715
# with at least one defective = 4845-715
-------------------------------------------
C) How many will contain 1 defective calculator?
7C1*13C3 = 7*286 = 2002
-------------------------------------
D) How may of will contain 3 defective calculators?
7C3*13C1 = 13*(7*6*5)/(1*2*3) = 455
----------------
Cheers,
Stan H.
-----------------